### v1

# On Fundamental Forms and Warped Products

Abstract.
Warped products are one of the simplest families of Riemannian manifolds that can have non-trivial geometries. In this letter, we characterize the geometry of hypersurface embeddings arising from warped product manifolds using the language of higher fundamental forms.
Keywords: Riemannian geometry, warped product, hypersurface embedding, higher fundamental form
2010 Mathematics Subject Classification:
53C18, 53A55, 53C21, 58J32.
Acknowledgements
SB is supported by the Operational Programme Research Development and Education Project No. CZ.02.01.01/00/22-010/0007541. JŠ is supported by the Czech Science Foundation (GACR) grant GA22-00091S.
1. Introduction
The notion of a warped product manifold was introduced by Bishop and O’Neill in 1969 [2]. Given two (pseudo-)Riemannian manifolds $(F,g_{B})$ and $(B,g_{B})$, a warped product manifold $F\times_{f}B$ is a manifold $F\times B$ equipped with a (pseudo-)Riemannian metric
(1.1) $g=f^{2}g_{F}+g_{B}\,,$
where $f\in C^{\infty}_{+}(B)$. Here, $B$ is called the base manifold and $F$ is called the fiber manifold. Such a structure is emblematic of some of the simplest non-trivial geometries, and as such finds broad application in general relativity. For example, both the Schwarzschild metric and the FLRW metric are warped product manifolds and are physically relevant solutions to the Einstein field equations. Indeed, partial differential equations are separable on these manifolds, and hence provide a fertile ground for solving partial differential equations in non-trivial settings.
In this letter, we use elementary methods to investigate the properties of hypersurface embeddings in the context of warped product manifolds. The primary tool we will use to study these embeddings are fundamental forms.
In 1867, Bonnet proved the fundamental theorem of surfaces, which, in modern language, asserts that up to rigid Euclidean motions, a surface embedded in $\mathbb{R}^{3}$ is uniquely determined by its first and second fundamental forms. This result easily generalizes to hypersurfaces embedded in $\mathbb{R}^{d}$. The naming of these tensors begs for explanation: is there a third fundamental form? A fourth? Indeed there are—but in the context of Euclidean embeddings, they either vanish or are easily expressible in terms of the first and second fundamental forms [1, Volume 3] (depending on the definition provided). However, recent work [3] found that higher fundamental forms in a different setting have interesting geometric interpretations. It is the goal of this article to show that these fundamental forms play a special role in characterizing warped product manifolds that arise from hypersurface embeddings.
To that end, consider the following definitions.
Definition 1.1.
Let $\iota:\Sigma\hookrightarrow(M,g)$ be a smooth hypersurface embedding and let $n\in\Gamma(TM)$ be any vector orthogonal $T\Sigma$ such that $g(n,n)|_{\Sigma}=1$. The first fundamental form $\bar{g}:=\iota^{*}g$ is the pullback of $g$ to the hypersurface; this is also called the induced metric. The second fundamental form is defined according to ${\rm II}:=\iota^{*}\nabla g(n,\cdot)$. We define the third fundamental form according to ${{{\bf\rm III}}{\hskip 0.56905pt}}{}:=\iota^{*}g(R(n,\cdot)n,\cdot)$, and finally, for $k\geqslant 4$, we define the $k$th fundamental form using abstract index notation:
$\underline{\overline{\rm{k}}}_{ab}:=\iota^{*}(n^{a_{1}}\cdots n^{a_{k-3}}n^{c}n^{d}\nabla_{a_{1}}\cdots\nabla_{a_{k-3}}R_{cabd})\,.$
For the definition of the symbols and notations used above and in what follows, see Subsection 1.1. As is shown in Lemma 2.1, this definition provides a natural extension of the second fundamental form, insofar as they are related to normal derivatives of $\nabla n$.
In Section 2, we first prove a series of simple results that relate the fundamental forms and various derivatives of the metric with respect to the transverse direction, culminating in Theorem 2.7 which generalizes the fundamental theorem of surfaces. In Section 3, we use that result to identify the structure of hypersurface embeddings that can be realized as a warped product manifold where one of the fibers is the hypersurface—see Theorem 3.1. Finally, in Section 4, we characterize when a hypersurface embedding is realizeable as the base of a warped product manifold—see Theorems 4.1 and 4.5.
1.1. Notations and Conventions
In this letter, we denote by the pair $(M,g)$ a smooth $d$-dimensional Riemannian manifold equipped with a metric $g$. On this manifold we denote by $\nabla$ the Levi-Civita connection with Riemann curvature tensor $R$, defined via the convention
$R(x,y)z=\nabla_{x}\nabla_{y}z-\nabla_{y}\nabla_{x}z-\nabla_{[x,y]}z\,.$
Here, we have that $x,y,z\in\Gamma(TM)$ are vector fields in $M$ and $[x,y]$ is the Lie derivative of $y$ with respect to $x$.
Now, for ease of computation, we use abstract index notation: we label sections of a tensor product of copies of the tangent bundle and cotangent bundle with Latin letters from the beginning of the alphabet (so $a$, $b$, $c$, etc.) to denote their tensor type. Then, a vector field $x$ would be labelled as $x^{a}$ and a covector (or a 1-form) $\omega$ would be labelled as $\omega_{a}$. With this notation, tensor contraction is indicated with repeated upper and lower indices. It follows that the above definition of the Riemann curvature tensor is equivalent to the definition
$R_{ab}{}^{c}{}_{d}z^{d}=\nabla_{a}\nabla_{b}z^{c}-\nabla_{b}\nabla_{a}z^{c}\,.$
For notational brevity, when it is useful, we will place the symbol for a (co)vector field in the subscript of a tensor to indicate contraction; for example, if $x$ and $y$ are vector fields and $T_{ab}$ is some tensor field, we may write $T_{xy}:=x^{a}y^{b}T_{ab}$.
In the context of hypersurface embeddings, we label the induced tensors, connection, and curvatures with an overline $\bar{\bullet}$; for example, we would write the induced Levi-Civita connection on a hypersurface by $\bar{\nabla}$, and the corresponding Riemann curvature tensor as $\bar{R}$. Also note that we will use the same indices for tensor structures on a hypersurface because of the canonical isomorphism between $TM|_{\Sigma}$ and $T\Sigma\oplus\langle n\rangle$, but when relevant we will use indices from the middle of the Latin alphabet ($i$, $j$, and $k$, for example) to denote hypersurface directions. We will also indicate that two expressions are equivalent along the hypersurface via the stacked equality =$\stackrel{{\scriptstyle\Sigma}}{{=}}$.
2. Hypersurface Embeddings and Fundamental Forms
In general, for any hypersurface embedding $\Sigma\hookrightarrow(M,g)$, we may express the metric (in a collar neighborhood) in the normal form, given by
$g=dt^{2}+\bar{g}(t,x^{k})_{ij}dx^{i}dx^{j}\,,$
where $\bar{g}(t_{0},x^{k})$ is the induced metric on the hypersurface defined by $t=t_{0}$ and coordinates $\{x^{i}\}$ are defined by parallel transporting coordinates on $\Sigma$ along geodesics generated by $\partial_{t}$. We shall call $\partial_{t}$ conormal vector; it provides a canonical extension of the unit normal along $\Sigma$ to $M$. This normal form is quite useful, as it allows us to easily establish alternative formulas for the fundamental forms. Indeed, we have the following technical lemmas.
Lemma 2.1.
Let $\iota:\Sigma\hookrightarrow(M,g)$ be a smooth embedded hypersurface with unit conormal vector $n$. Then, for every $k\geqslant 2$, we have that
$\underline{\overline{\rm{k}}}=\iota^{*}\nabla_{n}^{k-2}\nabla n+\text{productsof lower order fundamental forms.}$.
Proof.
We prove this lemma by induction. Clearly, when $k=2$, the identity is true by explicit computation. Now, suppose that the identity holds for all $2\leqslant n\leqslant k$. Now, we compute:
$\iota^{*}\nabla_{n}^{k-1}\nabla_{a}n_{b}=$ $\iota^{*}\nabla_{n}^{k-2}n^{c}\nabla_{a}\nabla_{c}n_{b}+\iota^{*}\nabla_{n}^{k-2}n^{c}R_{cabd}n^{d}$
$=$ $\iota^{*}\nabla_{n}^{k-2}\nabla_{a}\nabla_{n}n_{b}-\iota^{*}\nabla_{n}^{k-2}[(\nabla_{a}n^{c})(\nabla_{c}n_{b})]+\iota^{*}\nabla_{n}^{k-2}R_{nabn}$
$=$ $\iota^{*}\nabla_{n}^{k-2}R_{nabn}-\iota^{*}\nabla_{n}^{k-2}[(\nabla_{a}n^{c})(\nabla_{c}n_{b})]$
$=$ $\iota^{*}n^{c}n^{d}n^{a_{1}}\cdots n^{a_{k-2}}\nabla_{a_{1}}\cdots\nabla_{a_{k-2}}R_{cabd}-\iota^{*}\nabla_{n}^{k-2}[(\nabla_{a}n^{c})(\nabla_{c}n_{b})]$
$=$ $\underline{\overline{\rm{k+1}}}{}_{ab}-\iota^{*}\nabla_{n}^{k-2}[(\nabla_{a}n^{c})(\nabla_{c}n_{b})]\,.$
Note that the third and fourth identities follow because, from the existence of the normal form, there exists a function $t$ such that $n=dt$, and so
$\nabla_{n}n_{a}=n^{c}\nabla_{c}\nabla_{a}s=n^{c}\nabla_{a}n_{c}=\tfrac{1}{2}\nabla_{a}n^{2}=0\,.$
Now by the Leibniz property and the induction hypothesis, it follows that $\iota^{*}\nabla_{n}^{k-2}[(\nabla_{a}n^{c})(\nabla_{c}n_{b})]$ can be expressed along $\Sigma$ as a sum of products of lower order fundamental forms. The lemma follows. ∎
Remark 2.2.
Note that the definition of the third fundamental form given in [1, Volume 3] is $-\iota^{*}\nabla_{n}\nabla_{a}n_{b}$, and so Lemma 2.1 shows that Definition 1.1 is a reasonable extension to higher fundamental forms.
Using this result, we have another straightforward lemma. Here $\mathcal{L}$ denotes the Lie derivative.
Lemma 2.3.
Let $\iota:\Sigma\hookrightarrow(M,g)$ be a smooth embedded hypersurface with unit conormal vector $n$. Then, for every $k\geqslant 2$, we have that
$\underline{\overline{\rm{k}}}=\tfrac{1}{2}\iota^{*}\mathcal{L}_{n}^{k-1}g+\text{products of lower order fundamental forms.}$.
Proof.
We prove by induction. When $k=2$, the result follows by explicit computation. Next, suppose that the identity holds for all $2\leqslant n\leqslant k$. Then,
$\tfrac{1}{2}\iota^{*}\mathcal{L}_{n}^{k}g_{ab}=$ $\tfrac{1}{2}\iota^{*}\mathcal{L}_{n}\mathcal{L}_{n}^{k-1}g_{ab}$
$=$ $\tfrac{1}{2}\iota^{*}\nabla_{n}\mathcal{L}_{n}^{k-1}g_{ab}+\iota^{*}(\mathcal{L}^{k-1}_{n}g)_{c(a}(\nabla_{b)}n^{c})$
$=$ $\tfrac{1}{2}\iota^{*}\nabla_{n}\mathcal{L}_{n}^{k-1}g_{ab}+\underline{\overline{\rm{k}}}_{c(a}{\rm II}_{b)}^{c}+\text{products of lower order fundamental forms}$
$=$ $\iota^{*}\nabla_{n}^{k-1}\nabla_{a}n_{b}+\text{products of lower order fundamental forms.}$.
But from Lemma 2.1, the last line can be expressed in terms of $\underline{\overline{\rm{k+1}}}$, and so induction follows. ∎
Remark 2.4.
Note that in all of these constructions involving equality modulo lower order terms, we may ascribe weights to better control which fundamental forms appear in each identity. Indeed, assigning a weight of $3-k$ to $\underline{\overline{\rm{k}}}$ and a weight of $-2$ for every contraction, it is clear that the weights of each summand in the above identities are always equal.
Additionally, observe that when the lower fundamental forms vanish, the original definition of the fundamental forms is equivalent to the constructions provided in Lemmas 2.1 and 2.3.
Using Lemma 2.3, for any fundamental form, one may expres it as a sum of Lie derivatives of the metric and their contractions:
Corollary 2.5.
Let $\iota:\Sigma\hookrightarrow(M,g)$ be a smooth embedded hypersurface with unit conormal vector $n$. Then for every $2\leqslant k$, there exists a homogeneous formula for $\underline{\overline{\rm{k}}}$ given by
$\underline{\overline{\rm{k}}}=\tfrac{1}{2}\iota^{*}\mathcal{L}_{n}^{k-1}g+F_{k}(\bar{g}^{-1},\iota^{*}\mathcal{L}_{n}g,\ldots,\iota^{*}\mathcal{L}_{n}^{k-2}g)\,.$
These formulas can be explicitly computed for any choice of $k$ simply by a recursive algorithm.
Lemma 2.3 also implies that the metric in a sufficiently small neighborhood around a point in $\Sigma$ (where $\Sigma$ is analytically emebedded in $(M,g)$) is uniquely determined by the induced metric on $\Sigma$ and the infinite family of fundamental forms for the embedding:
Proposition 2.6.
Let $\Sigma\hookrightarrow(M,g)$ be an analytic hypersurface embedding. Then, for any point $p\in\Sigma$, there exists a sufficiently small neighborhood $U_{p}$ around $p$ such that $g|_{U}$ is uniquely determined by the induced metric and the infinite family of fundamental forms.
Proof.
Using Corollary 2.5, one may invert the formula for each $\underline{\overline{\rm{m+1}}}$ to find a formula for each Lie derivative of $\iota^{*}\mathcal{L}_{n}^{m}g$ in terms of the $\{\bar{g}^{-1},{\rm II},\ldots,\underline{\overline{\rm{m+1}}}\}$. Now, such an embedding may always be expressed in terms of a normal form, so that
$g=dt^{2}+\bar{g}(t,x^{k})_{ij}dx^{i}dx^{j}\,,$
where component functions $\bar{g}(t,x^{k})_{ij}$ are analytic. But then $\bar{g}(t,x^{k})$ is uniquely determined (in a sufficiently small neighborhood around a point $p\in\Sigma$) by its Taylor series expansion about $t=0$. But because $n=\partial_{t}$ and $\partial_{a}\partial_{t}=0$, it follows that $\iota^{*}\mathcal{L}_{n}^{m}g=(\partial_{t}^{m}\bar{g})(0,x^{k})_{ij}dx^{i}dx^{j}$, and so $\bar{g}(t,x^{k})$ is determined uniquely (in a sufficiently small neighborhood) by the family of tensors $\{\iota^{*}\mathcal{L}_{n}^{m}g\}_{m=0}^{\infty}$. However, as established, there exists a formula for each of these Lie derivatives in terms of $\bar{g}^{-1}$ and the fundamental forms. The proposotion follows. ∎
This uniqueness result is quite powerful, as it allows us to compare the fundamental forms of hypersurface embeddings to determine whether those hypersurface embeddings are equivalent.
Theorem 2.7.
Let $\iota:\Sigma\hookrightarrow(M,g)$ and $\hat{\iota}:\Sigma\hookrightarrow(M,\hat{g})$ be two analytic hypersurface embeddings such that, for every $p\in\Sigma$, $\iota(p)=\hat{\iota}(p)$. Then, in a collar neighborhood around $\Sigma$, $\hat{g}=g$ if and only if $\iota^{*}g=\hat{\iota}^{*}\hat{g}$ and, for every $k$, we have that $\underline{\overline{\rm{k}}}=\hat{\underline{\overline{\rm{k}}}}$.
Proof.
First, suppose that $\hat{g}=g$. Then it follows trivially that all geometric quantities are equal, including the induced metric and the fundamental forms. On the other hand, the other direction follows directly from the uniqueness of Proposition 2.6 and the existence of the normal form. ∎
As a consequence of Theorem 2.7, we can determine when a hypersurface embedding is actually a product manifold.
Corollary 2.8.
Let $\Sigma\hookrightarrow(M,g)$ be an analytic hypersurface embedding. Then $\underline{\overline{\rm{k}}}=0$ for every $k\geqslant 2$ if and only if in a collar neighborhood $U:=(-\epsilon,\epsilon)\times\Sigma$ of $\Sigma$, $(U,g|_{U})$ is a product manifold.
Proof.
We define a product manifold on $(-\epsilon,\epsilon)\times\Sigma$ by the metric $g=dt^{2}+\iota^{*}g$. Clearly, this product manifold has vanishing fundamental forms and the induced metric on $\Sigma$ agrees with that of $\Sigma\hookrightarrow(M,g)$. Hence by Theorem 2.7, these metrics agree. The other direction is trivial. ∎
3. Fiber-like Hypersurface Embeddings
Now, when a hypersurface $\Sigma\hookrightarrow(M,g)$ can be viewed as the fiber in a warped product, there exists a choice of local coordinates on $M$ such that the metric takes the form
$g=dt^{2}+h(t)\bar{g}_{ij}dx^{i}dx^{j}\,,$
where $\Sigma=\mathcal{Z}(t)$ is zero locus of $t$, coordinates $\{x^{i}\}$ are coordinates on $\Sigma$ parallel transported by $\partial_{t}$, $\bar{g}_{ij}$ depends only on $x^{i}$, and $h\in C^{\infty}_{+}(\mathbb{R})$. We call such a hypersurface embedding a fiber-like hypersurface embedding. Note this corresponds to (1.1) with $h=f^{2}$ and 1-dimensional base manifold.
We can then use the results of the previous section to determine when a hypersurface embedding is indeed a fiber-like hypersurface embedding. We thus have one of the main theorems of this letter.
Theorem 3.1.
Let $\Sigma\hookrightarrow(M,g)$ be an analytic hypersurface embedding with $\bar{g}:=\iota^{*}g$. Then, a collar neighborhood around $\Sigma$ is isometric to a fiber-like hypersurface embedding if and only if there exists analytic $h\in C^{\infty}_{+}(\mathbb{R})$ such that for every $k\geqslant 2$,
$\underline{\overline{\rm{k}}}=\tfrac{1}{2}h^{(k-1)}(0)\,\bar{g}+F_{k}(\bar{g}^{-1},h^{\prime}(0)\,\bar{g},\ldots,h^{(k-2)}(0)\,\bar{g})\,.$
Proof.
First, suppose that $\Sigma\hookrightarrow(M,g)$ is an analytic fiber-like hypersurface embedding, meaning there exists some analytic $h\in C^{\infty}_{+}(\mathbb{R})$ such that, in the normal form,
$g=dt^{2}+h(t)\bar{g}_{ij}dx^{i}dx^{j}\,.$
Clearly, because $n=\partial_{t}$, it follows that $\iota^{*}\mathcal{L}_{n}^{m}g=\iota^{*}(\partial_{t}^{m}h)\bar{g}$. But from Corollary 2.5, it follows that
$\underline{\overline{\rm{k}}}=\tfrac{1}{2}\iota^{*}\mathcal{L}_{n}^{k-1}g+F_{k}(\bar{g}^{-1},\iota^{*}\mathcal{L}_{n}g,\ldots,\iota^{*}\mathcal{L}_{n}^{k-2}g)=\tfrac{1}{2}h^{(k-1)}(0)\,\bar{g}+F_{k}(\bar{g}^{-1},h^{\prime}(0)\,\bar{g},\ldots,h^{(k-2)}(0)\,\bar{g})\,.$
This gives us one direction.
On the other hand, suppose that there exists an analytic $h\in C^{\infty}_{+}(\mathbb{R})$ such that for every $k\geqslant 2$,
$\underline{\overline{\rm{k}}}=\tfrac{1}{2}h^{(k-1)}(0)\,\bar{g}+F_{k}(\bar{g}^{-1},h^{\prime}(0)\,\bar{g},\ldots,h^{(k-2)}(0)\,\bar{g})\,.$
Now, the metric $g$ is expressible in normal form as $g=dt^{2}+\bar{g}(t,x^{k})_{ij}dx^{i}dx^{j}$. So, define a metric on a collar neighborhood of $\Sigma$ by $\hat{g}=dt^{2}+h(t)\bar{g}(0,x^{k})_{ij}dx^{i}dx^{j}$. However, from the above considerations, this metric $\hat{g}$ induces fundamental forms that agree exactly with those of the original embedding. Hence by Theorem 2.7, $\hat{g}=g$, and so $\Sigma\hookrightarrow(M,g)$ is isometric to a fiber-like hypersurface embedding in a collar neighborhood. ∎
Remark 3.2.
Explicitly, one may show that, in the language of the above theorem,
${{{\bf\rm III}}{\hskip 0.56905pt}}{}_{ab}=\tfrac{1}{2}h^{\prime\prime}(0)\bar{g}-\tfrac{1}{4}\bar{g}^{cd}(h^{\prime}(0)\bar{g}_{ac})(h^{\prime}(0)\bar{g}_{bd})\,.$
4. Base-like Hypersurface Embeddings
Warped product manifolds with a hypersurface embedding $\Sigma\hookrightarrow(M,g)$ as the base manifold have a much richer structure. In this case, there exists some function $f\in C^{\infty}_{+}(\Sigma)$ and some coordinates $(t,x^{i})$ such that
$g=f(x^{i})^{2}dt^{2}+\bar{g}_{ij}dx^{i}dx^{j}\,.$
Here $\Sigma=\mathcal{Z}(t)$ is zero locus of $t$, coordinates $\{x^{i}\}$ are coordinates on $\Sigma$ parallel transported by $\partial_{t}$, and $\bar{g}_{ij}$ depends only on $x^{i}$. We call these manifolds base-like hypersurface embeddings. Such embeddings can also be characterized in terms of fundamental forms, but their behavior is much more subtle. Note this corresponds to (1.1) with $(d-1)$-dimensional base manifold.
First, we prove that for any base-like hypersurface embedding, the even fundamental forms vanish.
Theorem 4.1.
Let $\Sigma\hookrightarrow(M,g)$ be a base-like hypersurface embedding with warping function $f$. Then, $\underline{\overline{\rm{k}}}=0$ for all $k=2n$ for positive integers $n$.
Proof.
As $(M,g)$ is a warped product, in a coordinate system adapted to the warped product $(t,x^{i})$, the metric is expressible as
$g=f(x^{1},\ldots,x^{d-1})^{2}dt^{2}+\bar{g}_{ij}(x^{1},\ldots,x^{d-1})dx^{i}dx^{j}\,.$
Now we wish to change coordinates so that $g$ is asymptotically in normal form, i.e. so that
$g=F^{(n+2)}(s,y^{j})ds^{2}+G_{i}^{(n+1)}(s,y^{j})dsdy^{i}+\gamma_{ij}(s,y^{1},\ldots,y^{d-1})dy^{i}dy^{j}\,,$
where $n\geqslant 0$ is even and $F^{(n+2)}(s,y^{j})=1+\mathcal{O}(s^{n+2})$ is an even function of $s$. Similarly, we will require that $G_{i}^{(n+1)}=\mathcal{O}(s^{n+1})$ is an odd function of $s$.
To find such a coordinate transformation, we solve order-by-order inductively. The base case is fairly easy, with $n=0$: define $t=\frac{s}{f(y^{i})}$ and $x^{i}=y^{i}$. Then, we have that
$g=ds^{2}-\frac{2s\partial_{y^{i}}f}{f}dy^{i}ds+\gamma_{ij}dy^{i}dy^{j}\,,$
for some $\gamma_{ij}$. Clearly, here $F^{(2)}(s,y^{j})=1$ is even and $G_{i}^{(1)}(s,y^{j})=-\frac{2s\partial_{y^{i}}f}{f}$ is odd in $s$. We now use induction to proceed.
Suppose that $n\geqslant 0$ is even, that there exists $T_{n+1}(s,y^{j})=\frac{s}{f(y^{j})}+\cdots+Q_{n+1}(y^{j})s^{n+1}$ an odd polynomial in $s$, and $X^{i}_{n}(s,y^{j})=y^{i}+\cdots+P^{i}_{n}(y^{j})s^{n}$ an even function of $s$ such that
$g=F^{(n+2)}(s,y^{j})ds^{2}+G_{i}^{(n+1)}(s,y^{j})dsdy^{i}+\gamma_{ij}dy^{i}dy^{j}$
when we use coordinates $t=T_{n+1}(s,y^{j})$ and $x^{i}=X^{i}_{n}(s,y^{j})$, where $F^{(n+2)}$ is an even function of $s$ (with first coefficient 1) and $G_{i}^{(n+1)}$ is an odd function of $s$. When $n=0$, this is the base case.
We now increment $n\rightarrow n+2$. Let $t=T_{n+1}(s,y^{j})+s^{n+3}A_{n+3}(y^{j})$ and $x^{i}=X^{i}_{n}(s,y^{j})+s^{n+2}B^{i}_{n+2}(y^{j})$, with $A_{n+3}$ and $B^{i}_{n+2}$ arbitrary functions of $\vec{y}$. We would like to show that there exist unique $A_{n+3}$ and $B^{i}_{n+2}$ such that $g_{ss}=F^{(n+4)}$ and $g_{si}=G^{(n+3)}_{i}$. To do so, we may compute components of the metric:
$g_{si}=$ $G^{(n+1)}_{i}+(n+2)s^{n+1}\gamma_{jk}(\partial_{y^{i}}X_{n}^{j})B^{k}_{n+2}$
$+s^{n+2}\left((n+3)g_{tt}A_{n+3}\partial_{y^{i}}T_{n+1}+h_{jk}(\partial_{s}X^{j}_{n})\partial_{y^{i}}B^{k}_{n+2}\right)$
$+s^{n+3}g_{tt}(\partial_{s}T_{n+1})\partial_{y^{i}}A_{n+3}$
$+(n+2)s^{2n+3}h_{jk}B^{j}_{n+2}\partial_{y^{i}}B^{k}_{n+2}+(n+3)s^{2n+5}g_{tt}A_{n+3}\partial_{y^{i}}A_{n+3}\,.$
Because $h_{jk}$ is invertible, we may fix $B^{i}_{n+2}$ so that the second term cancels with the leading term of $G^{(n+1)}_{i}$. Because $G^{(n+1)}_{i}$ is odd and $X^{j}_{n}$ is even (satisfying $\partial_{y^{i}}X^{j}_{n}=\delta_{i}^{j}+\mathcal{O}(s^{2})$), the remaining higher-order terms are shunted to order $n+3$ and remain odd. Then, note that because both $T_{n+1}$ and $\partial_{s}X^{j}_{n}$ are odd in $s$, we see that the second line is actually odd and is of order $s^{n+3}$. Indeed, it follows that $g_{si}$ is odd and of order $s^{n+3}$. Now having fixed $B^{i}_{n+2}$, we may consider $g_{ss}$:
$g_{ss}=$ $1+2(n+2)s^{n+1}h_{jk}B^{j}_{n+2}\partial_{s}X^{k}_{n}$
$+(F^{(n+2)}-1)+2(n+3)s^{n+2}g_{tt}A_{n+3}\partial_{s}T_{n+1}$
$+(n+2)^{2}s^{2n+2}h_{jk}B^{j}_{n+2}B^{k}_{n+2}+(n+3)^{2}s^{2n+4}g_{tt}A_{n+3}^{2}\,.$
Now because $B^{i}_{n+2}$ is fixed and $\partial_{s}X^{k}_{n}$ is odd, the second term on the first line is shunted to the second line (and remains even). Thus, noting that $\partial_{s}T_{n+1}|_{s=0}\neq 0$, we may choose $A_{n+3}$ uniquely so that the second line vanishes to leading order. But again, $\partial_{s}T_{n+1}$ is even as is $F^{(n+2)}$, so all higher-order terms remain even. Thus, the induction holds.
Given such a coordinate transformation, we can now determine the parity of $\gamma_{ij}$. However, this is trivially the case: $\frac{dt}{dy^{i}}$ is odd, so $\frac{dt}{dy^{i}}\frac{dt}{dy^{j}}$ is even. Similarly, $\frac{dx^{a}}{dy^{i}}\frac{dx^{b}}{dy^{j}}$ is even. Therefore, we find that $\gamma_{ij}$ is an even function of $s$. Taking stock of this coordinate transform, we have that
$g=ds^{2}+(\gamma^{(0)}_{ij}+s^{2}\gamma^{(2)}_{ij}+\cdots)dy^{i}dy^{j}+\mathcal{O}(s^{m})$
for any choice of $m$. Therefore, we have that in these coordinates, $n^{a}=\partial_{s}+\mathcal{O}(s^{m})$, and so $\mathcal{L}_{n}T=\partial_{s}T+\mathcal{O}(s^{m-1})$ for any tensor $T$. So, for any $k\leqslant m$, we have that $\iota^{*}\mathcal{L}_{n}^{k}g=\iota^{*}\partial_{s}^{k}g$. Notably, for any $k$ odd, we have that we may always find a coordinate transform so that
$\iota^{*}\partial_{s}^{k}g=0\,.$
Now recall from Lemma 2.3 that $\iota^{*}\mathcal{L}_{n}^{k}g$ is proportional to the $(k+1)$th fundamental form, modulo lower order fundamental forms. So it follows that when $k$ is odd,
k+1¯¯=lower order fundamental $\underline{\overline{\rm{k+1}}}=\text{lower order fundamental forms.}$.
Here, as per Remark 2.4, each summand on the right hand side of the above display has odd weight $2-k$. But odd fundamental forms always have even weight, as does $\bar{g}^{-1}$. It thus follows that no summand on the right hand side of the above display is composed entirely of odd fundamental forms. Hence, it follows that if $\underline{\overline{\rm{m}}}$ vanishes for each $12\leqslant m\leqslant k-$ 1 even, then it also follows that $\underline{\overline{\rm{k+1}}}$ vanishes. This completes the induction and hence the proof.
This theorem shows that the first potentially non-vanishing fundamental form for a base-like hypersurface embedding is ${{{\bf\rm III}}{\hskip 0.56905pt}}{}$. It follows from the definition of ${{{\bf\rm III}}{\hskip 0.56905pt}}{}$ and [4, Proposition 42] that
${{{\bf\rm III}}{\hskip 0.56905pt}}{}=f^{-1}\bar{\nabla}^{2}f\,.$
This observation is essential for the results that follow.
Proposition 4.2.
Let $\Sigma\hookrightarrow(M,g)$ be a base-like hypersurface embedding. Furthermore, let ${{{\bf\rm III}}{\hskip 0.56905pt}}{}=0$. Then, in a neighborhood around any point $p\in\Sigma$, $(M,g)$ is isometric to a product manifold.
Proof.
First, note that if ${{{\bf\rm III}}{\hskip 0.56905pt}}{}=0$, then ¯$\bar{\nabla}^{2}f=0$. It therefore follows that both ¯$\bar{\nabla}f$ and $\iota^{*}g-df\otimes df$ are parallel with respect to $\bar{\nabla}$. Therefore, for any point $p\in\Sigma$, there exists a neighborhood $\bar{U}\subset\Sigma$ around $p$ with coordinates $(f,y^{m})$ such that the induced metric satisfies
$\iota^{*}g|_{\bar{U}}=df^{2}+\bar{\gamma}_{mn}dy^{m}dy^{n}\,.$
Now, by parallel transporting these coordinates transverse to $\Sigma$, these coordinates are defined on a small neighborhood $U\subset M$ around $p$ such that the metric becomes
$g|_{U}=f^{2}dt^{2}+df^{2}+\bar{\gamma}_{mn}dy^{m}dy^{n}\,.$
But then the standard coordinate transformation to cartesian coordinates $f=\sqrt{u^{2}+v^{2}}$ and $t=\arctan\tfrac{v}{u}$ yields a product metric (at least locally):
$g=dv^{2}+du^{2}+\bar{\gamma}_{mn}dy^{m}dy^{n}\,,$
and here $\Sigma=\mathcal{Z}(v)$ is the zero locus of $v$. The proposition follows. ∎
But from this proposition and from Corollary 2.8, it follows that each fundamental form vanishes, which suggests that the odd fundamental forms must be expressible as homogeneous operators on ${{{\bf\rm III}}{\hskip 0.56905pt}}{}$. Indeed, we have the following proposition:
Proposition 4.3.
Let $\Sigma\hookrightarrow(M,g)$ be a base-like hypersurface. Then, there exists an infinite family $\{O^{f}_{n}\}_{n=1}^{\infty}$ of differential operators on $\Gamma(\odot^{2}T^{*}\Sigma)$ depending on $f$ such that
$\underline{\overline{\rm{2n+3}}}=O^{f}_{n}({{{\bf\rm III}}{\hskip 0.56905pt}}{})\,.$
Furthermore, this operator satisfies $O^{f}_{n}(0)=0$.
Proof.
First, note that $\underline{\overline{\rm{2n+3}}}$ has a local, differential formula expressible entirely in terms of $f$, $\bar{g}$, their partial derivatives, and $\bar{g}^{-1}$. Now, as such objects are tensorial, we can instead write such objects using formulas constructed entirely from $f$, its (hypersurface) covariant derivatives, the metric $\bar{g}$ and its inverse, and the curvature tensor $\bar{R}$.
Now in the case where $f$ is constant, we trivially have from Corollary 2.8 and Proposition 4.2 that $\underline{\overline{\rm{2n+3}}}=0$. Therefore, any such formula for $\underline{\overline{\rm{2n+3}}}$ cannot contain terms that only depend on $\bar{R}$ and $f$ — i.e. all summands in such a formula must contain one or more derivatives of $f$. Thus, we may decompose any such formula into the following pieces:
$\underline{\overline{\rm{2n+3}}}=O(\bar{\nabla}^{2}f)+P(\bar{\nabla}f)\,,$
where $O$ is some homogeneous (with respect to rescaling $f$ by a constant) differential operator and $P$ is a (potentially non-linear) homogeneous algebraic operator composed of the curvature tensor $\bar{R}$, its derivatives, the metric, its inverse, and $f$.
Now consider the case where ${{{\bf\rm III}}{\hskip 0.56905pt}}{}=0$, and thus ¯$\bar{\nabla}^{2}f=0$. Because $\underline{\overline{\rm{2n+3}}}=0$, it therefore follows that
$P(\bar{\nabla}f)=0\,.$
However, ¯$\bar{\nabla}f$ does not necessarily vanish (as it is independent of ¯$\bar{\nabla}^{2}f$). But because $P$ is an algebraic operator that vanishes for any choice of $f$ satisfying ¯$\bar{\nabla}^{2}f=0$, we must have that $P=0$ identically. It thus follows that $\underline{\overline{\rm{2n+3}}}=O(\bar{\nabla}^{2}f)=:O^{f}_{n}({{{\bf\rm III}}{\hskip 0.56905pt}}{})$. As $O$ is homogeneous, it trivially follows that $O^{f}_{n}(0)=0$. ∎
Remark 4.4.
As a first example, one may show, by direct computation, that for base-like hypersurface embeddings,
${{\underline{\overline{\bf\rm V}}}}{}_{ab}=f^{-1}(\bar{\nabla}^{c}f)(-3\bar{\nabla}_{c}{{{\bf\rm III}}{\hskip 0.56905pt}}{}_{ab}+2\bar{\nabla}_{(a}{{{\bf\rm III}}{\hskip 0.56905pt}}{}_{b)c})\,.$
Surprisingly, the operator $O^{f}_{1}$ is quite simple.
From Proposition 4.3, we finally have the following theorem.
Theorem 4.5.
Let $\Sigma\hookrightarrow(M,g)$ be an analytic hypersurface embedding. Further, suppose that $\underline{\overline{\rm{2n}}}=0$ for each $n$, that there exists a positive analytic function $f:\Sigma\hookrightarrow\mathbb{R}$ satisfying ¯$\bar{\nabla}^{2}f=f{{{\bf\rm III}}{\hskip 0.56905pt}}{}$, and that $\underline{\overline{\rm{2n+3}}}=O^{f}_{n}({{{\bf\rm III}}{\hskip 0.56905pt}}{})$ for each $n$. Then, in a neighborhood $U$ of a point in $\Sigma$, $(U,g|_{U})$ is isometric to a warped product manifold with base manifold $U\cap\Sigma$.
Proof.
Given the function $f$ and following Theorem 4.1 and Proposition 4.3, we may construct a base-like hypersurface embedding with induced metric and fundamental forms that match those given in the theorem. Hence it follows from Theorem 2.7, that, in a neighborhood of a point in $\Sigma$, we have that the given embedding restricts to a base-like hypersurface embedding.